OpenStudy (anonymous):

Solve the system of equations below by graphing them with a pencil and paper. Choose the correct ordered pair (b, u). u = -b + 21 u = -2b + 30

OpenStudy (anonymous):

-b+21=-2b+30 b=9 sub to u=-b+21 u=11 (9,11)

OpenStudy (anonymous):

what a nice picture http://www.wolframalpha.com/input/?i=y+%3D+-x+%2B+21%2C+y+%3D+-2x+%2B+30

OpenStudy (anonymous):

i wrong u=-9+21 it's -12 so (9,12)

OpenStudy (anonymous):

solution is at the bottom below the nice picture